OpenStudy (anonymous):

Let f(x)=|2x+10| and g(x)=|2x-24| . Find all values of x for which f(x)=g(x) .

OpenStudy (turingtest):

f(x)=g(x) |2x+10|=|2x-24| Case 1: 2x+10=2x-24 0=-34 No Solutions Case 2: 2x+10=-2x+24 4x=14 x=7/2 The other two cases are equivalent so the only solution is x=7/2 think about why the other two cases are equivalent.

OpenStudy (anonymous):

because one this time the opposite side is negative and when both sides become negative its like them being positive